a father is twice as old as his son. 20 years ago, the age of the father was 12 times the age of the son. the present age of the father is.
Answer by mananth(16946)

let son be x years
father = 2x
..
20 years ago
son = x-20
father = 2x-20
..
2x-20=12*(x-20)
2x-20 = 12x -240
add 20 to both sides
2x = 12x-240+20
2x=12x-220
add -12x to both sides
2x-12x= 12x-12x-220
-12x+2x= -240+20
-10x= -220
divide by -10
x= -220/-10
x=22 son's age
Father's age = 2*22 = 44 years